OpenStudy (mendicant_bias):

(Piecewise-defined Functions) If I have a piecewise function where both intervals are open and right next to each other, is the overlapping point defined? (Example below)

OpenStudy (mendicant_bias):

E.g. if I have a function f(x) equal to \[1, \ \ \ -\pi/2<x<\pi/2,\]\[0, \ \ \ \pi/2<x<3\pi/2.\] Is my function defined at pi/2? It isn't right?

OpenStudy (welshfella):

it isn;t correctft

OpenStudy (welshfella):

* correct

OpenStudy (phi):

If your function were defined at pi/2 , then you could say its value. But we can't tell the function's value at pi/2 ... so it's not defined

OpenStudy (mendicant_bias):

Could you elaborate? What you said meant literally nothing to me. "It isn't correct". What? What isn't correct? What I said, the problem itself? Thanks, Phi.

OpenStudy (mendicant_bias):

My question was stemming from something dealing w/Fourier Series, are you at all good with Mathematica?

OpenStudy (phi):

not really, but I can look at your question